OpenStudy (wormynut):

please help! i have my finals tommorow!! Sam scored 75,82 and 90 in 3 maths tests.How many marks must he obtain in th next test to have an average of exactly 85 for the 4 maths tests?

OpenStudy (welshfella):

let x be the score in the nest test then (75 + 82 + 90 + x) / 4 = 85 247 + x = 4*85 find x

OpenStudy (wormynut):

thankyoouuuuu!!